a) Find the indefinite integral of the function

𝑦 = 3𝑑2 + 2𝑒3𝑑 + 2 + 2 cos 3𝑑


1
Expert's answer
2022-02-15T17:51:26-0500
"\int ydt=\int (3t^2+2e^{3t} +2+2\cos 3t)dt"

"=t^3+\dfrac{2}{3}e^{3t}+2t+\dfrac{2}{3}\sin 3t+C"